Grade 12Algebra

(-3)^1/2 ×(-3)^1/2 =? What is the answer of the above question

madhu
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
As we realize that the result of two complex numbers is can yield a mind boggling number or Real number. E.g. (iy)* (ib) = - yb(a Real number). Like savvy we have (- 3)^1/2 ×(- 3)^1/2 we get √3 i* √3 i = - 3